OpenStudy (anonymous):

what is the inverse of e^(x+2)

OpenStudy (mathmagician):

if you have y=exp(x+2), then ln(exp(x+2))=ln(y) x+2=ln(y) x=ln(y)-2 inverse of y=f(x) is x=f(y), so x=ln(y)-2